### v2

# Beyond the EPICS: comprehensive Python IOC development with QueueIOC

Abstract
Architectural deficiencies in EPICS lead to inefficiency in the development and application of EPICS input/output controllers (IOCs). An unintrusive solution is replacing EPICS IOCs with more maintainable and flexible Python IOCs, only reusing the Channel Access (CA) protocol of EPICS. After a digression about GUI development inspired by EPICS operator interfaces (OPIs), the structural similarity between standalone GUI backends, the Mamba backend, EPICS IOCs and other server programs is analysed. By combining the caproto library and event loops like those in these server programs, the QueueIOC framework for Python IOCs is created, which has the potential to systematically replace most EPICS IOCs currently used. Examples are first given for workalikes of StreamDevice and asyn; examples for seq-like applications include monochromators, motor anti-bumping and motor multiplexing. Also shown is software to use with the ~/iocBoot convention which addresses some issues with a similar solution based on procServ, along with a workalike of procServControl. A QueueIOC-based framework for detector integration, which overcomes areaDetector’s limitations in performance and architecture, is presented in an accompanying paper.
1 Introduction
The Experimental Physics and Industrial Control System (EPICS) is a basis for accelerator control and beamline control at HEPS, the High Energy Photon Source (Chu et al., 2018; Liu, Dong and Li, 2022). The rich selection of EPICS modules available has greatly facilitated device-control tasks at HEPS. However, during these tasks, it also became apparent that EPICS has some deficiencies inherent in its architecture (cf. Figure 4b), which lead to inefficiency in development and application.
Our first complaints with EPICS are about its conception of “record links” (Kraimer et al., 2018), as well as the locking, scanning and processing mechanism associated with them. EPICS record links have different types and attributes: there are input links, output links and forward links; they can be subdivided into database links, Channel Access links and other subtypes; the “Process Passive” attribute can have the values NPP, PP, CA, CP and CPP. Events (hardware interrupts, user writing, periodic “scanning”) result in record “processing”, where the selection and ordering of records to be processed are determined by these links, according to the latter’s types and attributes. The rules for this selection and ordering are complex, and a major factor contributing to the complexity is the tight coupling between the types and attributes of record links. Ideally there should be separate rules about types, subtypes and attributes (especially “Process Passive” among the last) that add up to form the rules for individual combinations; the actual rules are quite far from able to be decomposed (the results would surely be easier to remember), and are instead quite like a lot of special cases piled together. The complexity of record links leads to considerable difficulty in learning: even among experienced EPICS application developers, few can precisely retell the rule for a less common combination of type and attribute. Fortunately, in practice only a handful of combinations are commonly used; on the other hand, even with the already complex set of rules, there are still requirements that cannot be naturally done, and a prime example for this is the motor module. In addition to the types and attributes of links, the selection and ordering of records to be processed are also affected by the records themselves, depending on them being “synchronous” or “asynchronous”; however, motor records are neither synchronous nor asynchronous.
Apart from the link mechanism, EPICS records themselves are also quite inexpressive: EPICS “databases”, which are collections of records, can be nested; however, nested database files cannot be loaded at runtime, and instead must be expanded at compile time; .substitutions files can be loaded at runtime, but it cannot be nested either. This makes it very hard to abstract reusable functionalities as pluggable sub-databases, even for simple requirements like proportional-integral-derivative (PID) controllers. One idiomatic alternative is writing these functionalities as customised records, eg. EPID from the std module, which involves a considerable amount of boilerplate code and quite an extent of knowledge on EPICS internals. Another common alternative is writing “sequencers” based on the seq module, which is often error-prone because these “sequencers” can easily become comparable to C/C++ code riddled with goto statements; the monochromator code in the optics module are representative examples for this issue. The st.cmd language in EPICS does not have looping or conditional constructs, so we also cannot circumvent the inexpressiveness of records on the st.cmd layer. To work around the inexpressiveness of st.cmd itself, it is common practice to use external code generators, eg. iocbuilder (Abbott and Cobb, 2011), makeDb.py (from the ADGenICam module) and our own ones (cf. ADXspress3 in the next paragraph). We note that as can be seen from the examples given, code generators are in particular commonly used to deal with hardware that expose dynamic interfaces: variable numbers of detector channels, variable sets of camera features, etc. st.cmd also feels inconvenient to operators, as it is mainly intended as a developer-oriented language, and its inexpressiveness further impedes user-oriented abstraction.
It is undoubted that none of the issues above is fatal to EPICS: developers can still produce usable, reliable EPICS modules, and users can still use these modules to actually fulfill their needs; as has been summarised above, the problems are about inexpressiveness and inconvenience. Meanwhile, during the construction of HEPS, we have found that improvements on expressiveness and convenience can be of great value in boosting efficiency; here we briefly give a few examples for this (Figure 1). The ihep-pkg packaging framework (Liu, Dong and Li, 2022, cf. also Section 4) now covers a large range of EPICS modules than that of the NSLS-II epicsdeb repository (Brookhaven National Laboratory, 2018), and provides support for CentOS 7, Rocky 8 plus partial support for Windows (using the MinGW environment), while still being easy to keep up to date. The ADXspress3 module (https://codeberg.org/CasperVector/ADXspress3) merges functionalities from both of its upstream versions (https://github.com/quantumdetectors/xspress3-epics and https://github.com/epics-modules/xspress3), and is much easier to maintain than both upstreams; meanwhile, to users it is also more convenient to build and use. The MambaPlanner mechanism (Li et al., 2023) provides a succinct, consistent command-line interface (CLI) for both step scans and fly scans, encapsulating hairy details about hardware configuration, correctness checks and data processing; while being convenient for debugging and advanced usage, it also facilitates development of graphical user interfaces (GUIs) oriented toward regular users. As is shown in Figure 1, they have each approached some kind of complexity lower-bound, in the sense that there is no obvious way to significantly simplify the code without inordinately complicating other code modules or sacrificing code clarity. Following the “succinctness is power” statement (Graham, 2002), by approaching the complexity lower-bound we can maximise the engineering efficiency. The examples above are just natural results of this rule; the efficiency boosts in them can be in 1–2 orders of magnitude, which also nicely coincide with the reduction of code needed to do the same tasks. Similarly, in other applications at HEPS, improvements in succinctness also prove to coincide with improvements in efficiency, even when the complexity lower-bounds have not yet been approached.
Figure 1: Some code examples, with the essential information in bold.
From the perspective above, we can give a summary of the problems with EPICS: the essence of these problems is that its architecture makes the complexities of EPICS modules vastly higher than the lower bounds; the latter can often be estimated intuitively, just like in Figure 1. We understand that the design of EPICS is deeply affected by historical factors: the choice of record and links over a full-fledged language, as well as the inexpressiveness of st.cmd, may be largely seen as results of hardware limitations on Versa Module Europa (VME) and similar hardware platforms; record and links are also reminiscent of the programming paradigm common with programmable logical controllers (PLCs), PandABox (Christian et al., 2019) etc. Nevertheless, with the fast growth of EPICS usage on PCs, it is also time to rethink about these design decisions. Of course, an option for completely new facilities is to embrace ecosystems with less historical burdens, eg. Karabo (Göries et al., 2023); for facilities with more legacies, a more backward-compatible option is to reuse the Channel Access (CA) protocol of EPICS, but internally use more succinct tools. Given the common programming practices in last decade (especially those related to large scientific facilities), Python is an obvious candidate for this task; in particular, we have chosen the pure-Python caproto library (Allan, 2021), as it provides succinct interfaces for both server-side and client-side programming with the CA protocol. However, the officially expected usage of caproto depends on Python’s async/await programming paradigm, which leads to additional costs in learning and development; in this paper, we describe how we solve this problem, and produce a general-purpose workalike of EPICS based on Python.
2 The submit/notify pattern for GUI programming
As has been mentioned above, succinctness and clarity are pursued in many fields of programming at HEPS; GUI programming is no exception from this. In this section, we will discuss a GUI programming pattern inspired by EPICS, which in turn helped to inspire our Python-based EPICS workalike besides being useful in itself. In our summary, the complexity of GUI programming has a few common sources, the first of which is the mixing of business logic with operations on widgets that represent the logic. This kind of mix-up is a quite well-known type of non-modularity in software engineering; it is generally resolved by decoupling GUI programs into frontends and backends, eg. those in Mamba (Liu et al., 2022; Dong et al., 2022). Another source of complexity is the complex interactions between widgets in GUI frontends, which complicates the call chains between these widgets. In most GUI frontends, there are some kind of quite complex states implicitly shared between widgets, which in conjunction with the strong concurrency usually associated with GUIs result in a significant difficulty in understanding transitions of the implicit states. This can easily result in timing problems, most often occurring as race conditions; it is yet another source of complexity, especially when coupled with complicated call chains.
GUIs in the EPICS ecosystem, known as operator interfaces (OPIs), can be easily composed in OPI editors in a drag-and-drop fashion. As drag and drop are inefficient in the creation of large OPIs with many repetitions, we are also systematically exploring PyDM (Slepicka, 2018), the Python-based OPI engine for this task. Notwithstanding this issue, we still find GUIs an advantage of EPICS, since we do not need to worry about the problems in the previous paragraph: OPI widgets do not directly interact or share states with each other, eliminating problems about call chains and timing. This is because the control logic in EPICS is gathered in input-output controllers (IOCs), and widgets are simply graphical representations of process variables (PVs); the latter are provided by IOCs, and manipulated by the control logic therein. Our attempts to systematically simplify the control logic in IOCs, by means of caproto-based Python IOCs, will be covered in the following sections, and here we continue our discussion on GUIs. Learning from OPIs, even in a general GUI we can require the widgets to communicate only with the main event loop (Figure 2a); to minimise state sharing, message passing can be mandated for this communication. To perform message passing, libraries like ZeroMQ and Python’s queue can surely be used; the signal/slot mechanism provided by some GUI libraries, eg. Qt and GTK, can also be readily used; here we additionally note that the CA protocol can also be regarded as a kind of message-passing mechanism.
Figure 2: (a) The submit/notify pattern, in comparison with (b) the MVC pattern; the latter image is courtesy of the Wikipedia entry “Model-view-controller”.
We call the design pattern in Figure 2(a) the submit/notify pattern, where widgets submit update requests (eg. setpoint inputs from the user) to the main event loop, and the main loop notifies widgets of actual updates (eg. readback values and other status changes). The interactions between a widget and the rest of the program are limited to the submissions it sends and the notifications it receives; the similar can be said for the main loop. In this way, the widgets and the main loop can be considered “actors” well decoupled from each other, as in the actor model of concurrent programming, which makes the program easy to reason about. The actor model, as well as its cousin model, communicating sequential processes (CSP), are highly influential: this can be seen in languages like Erlang and Go, as well as in libraries like MPI and ZeroMQ. Historically the model was also closely related to Smalltalk, the language well known for pioneering in object-oriented programming (OOP). If we also take the backend into consideration, the submit/notify pattern may also be compared with the model-view-controller (MVC) pattern (Figure 2b). The “view” part obviously corresponds to the widgets, and the backend is the real “model” part; the main loop should be a thin encapsulation of the backend, and thus also corresponds partially to the “model” part. The “controller” part is the submit/notify logic separated into the widgets’ event callbacks (or slot functions in the signal/slot mechanism) and the main loop.
As is hinted above, the main event loop in a GUI frontend should only contain logic and states essential to the graphical representation of the business logic – which in turn belongs to the backend; this improves the reliability and maintainability of GUIs, just like how privilege separation improves the security of qmail and software with similar designs (Bernstein, 2007). In the Mamba framework, the backend communicates with the frontends through a remote-procedure call (RPC) mechanism (Figure 2a) based on ZeroMQ. For tasks (eg. some in beamline control) that do not need the full Mamba infrastructure or even barebone Bluesky (Allan et al., 2019), we also developed some standalone GUIs with the submit/notify pattern; in these GUIs, there is still a clear separation between frontends and backends. Mamba frontends developed in the submit/notify style are, as of now, primarily used in our framework for automated attitude tuning of beamlines, presented in an accompanying paper (Li et al., 2024). The code for these frontends has been released in the open-source edition of Mamba (cf. Section 4); released in the same repository are two PyQt-based examples for standalone submit/notify GUIs. One of them is a workalike of the program in (Du et al., 2012), essentially an X-ray beam position monitor (XBPM) based on area detectors, with a simplified RPC mechanism between the backend and the frontend based on Python’s queue. The another is a simplified workalike of the ImageJ plugin provided by the ADViewers module, treating EPICS IOCs for area detectors as its backends, reusing the CA protocol for message passing with them.
3 The client-server model and EPICS IOCs
As has been discussed in Section 2, the CA protocol can be seen as a message-passing mechanism between servers (EPICS IOCs) and clients (EPICS OPIs etc); moreover, IOCs may be compared to GUI backends (whether standalone or in Mamba), and sometimes even be directly used as the backends. Therefore from analysing what is common in these servers, it is possible to find better ways to do what EPICS IOCs do; historically our standalone backends also learned from the Mamba backend, so here we begin with Mamba. Communication between the backend and frontends in Mamba may be categorised into requests/replies and notifications, the latter necessary due to the intrinsic weaknesses of state polling (Liu et al., 2022). In both standalone and Mamba backends there are main event loops: in Mamba backends, event handling (also including sending notifications) is separated into handlers registered to the core library, which encapsulates the main event loop; in standalone backends, there are explicit main loops that reply to requests and send notifications. We find the combination of an event loop, requests (with or without replies; an example for the latter is submissions in the submit/notify pattern, cf. Section 2) and notifications a common pattern in server programs. In addition to GUI backends, other examples are also abundant, eg. the alsamixer program to control audio volume under Linux (Figure 3): although not written with a frontend and a backend decoupled from each other, this program still features an event loop, which handles mouse/keyboard events and forwards status updates from the operating system. We are also aware of least one proprietary device-control product used at HEPS that explicitly uses requests/replies and notifications over TCP/IP network for the vendor’s application programming interface (API).
Figure 3: The alsamixer program.
In the light of these server programs, it is easy to find that caput, caget and camonitor, the basic operations in the CA protocol of EPICS, are also specialised combinations of requests/replies and notifications. Given these observations, we designed the basic architecture of our Python IOC framework, as in Figure 4(a). To accommodate the more mainstream programming style both inside and outside the Python ecosystem, the main event loop in the framework is currently based on a regular (non-async) function running in a regular Python thread. Correspondingly, there is a thin layer that isolates from the main loop the async/await-based code caproto expects (cf. Section 1); as the layer is based on variants of Python’s queue, we call our framework QueueIOC. For many IOCs in our framework the main loops are explicit, and are quite like the main loop in the standalone XBPM program in Section 2; however, for certain types of requirements exhibiting obvious regularity (eg. the QScanIOC-based IOCs in Section 4 and all IOCs in Section 5), we also encapsulate the boilerplate code in their main loops, so that the developer usually only needs to care about the essentials (cf. Figure 1, 5 & 8).
Figure 4: (a) The architecture of QueueIOC, in comparison with (b) the architecture of an EPICS IOC; the latter image is courtesy of Kraimer et al. (2018).
Aside from QueueIOC and barebone caproto, we are also aware of similar attempts at using Python in the EPICS ecosystem, eg. PyDevice (Vodopivec, 2020) and pythonSoftIOC (Cobb, 2021). Instead of comparing them in detail, here we note the most crucial difference between QueueIOC and the rest is that it attempts to replace most EPICS IOCs in a systematic way, while striving to keep the amounts of code for the Python IOCs satisfactorily close to their intuitive complexity lower-bounds. For the latter goal, we explicitly note that QueueIOC focuses on fulfilling needs in succinct ways, which are not necessarily how EPICS fulfills the same needs. In the following sections, concrete examples will be given for various types of applications doable with QueueIOC; we believe these examples can adequately show the potential of QueueIOC. We also note that motor IOCs are perhaps a most obvious type of IOCs currently not covered by QueueIOC, but that we do not find this a fundamental weakness. This is because the implementation of the motor record contains a state machine with a few thousands lines of code, as well as a similar amount of ancillary code, which would take considerable human resource to port to Python even if we omitted some less useful features. A simple motor-like interface is indeed provided in QueueIOC, but this interface is, at least for now, intended only for “sequencers” (cf. Section 5) and not real motors.
To fully understand the potential of QueueIOC, it is instructive to compare its architecture with the architecture of EPICS IOCs (Figure 4b). Database access, IOC database, record support and device support are either implicit or unnecessary in QueueIOC; device drivers and interrupt routines are implicit in the underlying control layer; monitors are implicit in the CA layer. By treating other IOCs (communicating through the CA protocol, cf. also Section 5) as the underlying control layer, “sequencers” are just a specialised kind of IOCs. With mechanisms like QScanIOC (cf. Section 4) and QSlowIOC (Zhang et al., 2024), what scanners do can also be done succinctly with QueueIOC. In summary, all architectural elements in EPICS IOCs have satisfactory counterparts in QueueIOC; so in the long term, we believe QueueIOC has the potential to be eventually capable of doing what is done with all EPICS IOCs currently used, except those IOCs that run under special resource constraints (most prominently VME IOCs). Furthermore, as has been noted above, caput, caget and camonitor are just specialised requests/replies and notifications; we think similar conclusions could be made about the PV Access (PVA) protocol added to EPICS in recent years. In conjunction with our observations about server programs, it is also a reasonable guess that similar conclusions could be made about the communication protocols in other device-control ecosystems, eg. TANGO and Karabo. If these guesses were sufficiently accurate, QueueIOC might become a starting point for some kind of unified device-control ecosystem, since QueueIOC intentionally encourages a programming style agnostic of the device-control ecosystem: the developer usually only needs to care about requests/replies and notifications – not caput, caget or camonitor.
4 Device IOCs and “soft” IOCs with QueueIOC
QueueIOC and s6-epics (the latter to be covered below) have been released, respectively, at https://codeberg.org/CasperVector/queue_iocs and https://codeberg.org/CasperVector/s6-epics; fully open-source editions of Mamba and ihep-pkg have been released, respectively, at https://codeberg.org/CasperVector/mamba-ose and https://codeberg.org/CasperVector/ihep-pkg-ose. QueueIOC itself depends on some patches for caproto; GUIs in the open-source edition of Mamba (including the standalone ones, which are in the mamba_lite subdirectory) depend on some patches for pyqtgraph; the supplementary materials include OPIs which depend on some patches for PyDM; all these currently HEPS-specific patches are available from the open-source edition of ihep-pkg.
Our actual introduction to QueueIOC begins with a workalike of StreamDevice, which is in our eyes a simplest yet most useful EPICS module: the QScanIOC class, in combination with classes like TimedRWPair; an example IOC for this, qscan_b2985, is given for the Keysight B2985 electrometer, shown in Figure 5. As can be seen from the figure, QScanIOC provides friendly encapsulation for periodically polling (“scanning”) of devices, while TimedRWPair (just like the asyn module) encapsulates communication via serial ports or TCP/UDP. Unlike StreamDevice which uses “protocol files”, essentially a still quite limiting domain-specific language (DSL), QScanIOC-based IOCs have native access to Python’s capability to process textual and/or binary data. As these IOCs also have full access to Python’s expressiveness, we can easily abstract boilerplate code, eg. with cmd_map in qscan_b2985, which would be clumsy to do with EPICS databases. Based on the extensive support for communication protocols available with Python (whether in standard libraries or third-party libraries), with QueueIOC it is also easy to write IOCs based on more complex interfaces like JSON, HTTP, modbus etc: eg. the qdet_eiger IOC for the Eiger detector uses the vendor’s protocols based on JSON/HTTP and ZeroMQ. qdet_eiger is based on QDetectorIOC, a QueueIOC-based framework for detector integration presented in an accompanying paper (Zhang et al., 2024), which overcomes the areaDetector framework’s limitations in performance and architecture. Also demonstrated therein is how QueueIOC works with C/C++ libraries from vendors.
Figure 5: Source code for the qscan_b2985 IOC.
QScanIOC is intended for devices with basically stateless interfaces, whose main event loops only differ in the logic that can be separated into on_init(), on_scan() and on_req(). This is why the main loop in QScanIOC does not need to be explicitly written for each device; for most other devices, explicit main loops are necessary. For training purposes in writing main loops, the qdet_nct IOC for the Tsuji (N)CT counter and the qdet_o974 IOC for the Ortec 974 counter are good examples, both based on QDetectorIOC. QScanIOC is not used as we want them to support areaDetector-like acquisition of a fixed (and possibly infinite) number of frames, a feature unfit for stateless implementations. qdet_nct can automatically detect the number of counter channels on the specified device, and adjust its set of PVs accordingly, without relying on code generators (cf. Section 1). QueueIOC’s automatic adaption to dynamic hardware interfaces (cf. also Figure 8) is based on Python’s type() mechanism for runtime creation of classes; based on this mechanism, we can do advanced abstractions similar to those in the ADGenICam and ADEiger IOCs with much more succinct code (Zhang et al., 2024). As can be seen from Figure 5, IOCs based on QueueIOC, whether for devices with dynamic interfaces or not, are easy to run for operators: unlike st.cmd in EPICS, with these Python IOCs device identifiers (addresses, ports etc) and other tuning parameters can be given on the command line, which is well separated from the IOCs’ source code. Additionally, when necessary it is also quite easy to customise IOC behaviours on deeper levels, without needing to copy entire IOC source files: eg. the file helpers/st_b2985.py shows how to customise the behaviours of qscan_b2985 in two aspects.
Till now all IOCs discussed in this section interact with hardware other than the controlling computer; besides “sequencers” which will be covered in Section 5, pure-software IOCs can also be written with QueueIOC. Our example for pure-software QueueIOC-based IOCs is qioc_s6, a workalike for the procServControl IOC, based on s6-epics, a workalike of procServ (Thompson, 2004). s6-epics itself is based on s6 (https://skarnet.org/software/s6/), a well-designed suite of programs to manage service processes; that latter most importantly (to us and concerning EPICS) supports separate logging for each service with reliable log rotation to avoid exhaustion of disk space. After organising IOCs in the ~/iocBoot convention, we can use administration commands like those in (Liu, Dong and Li, 2022), and a few additional useful commands (Figure 6). As can be seen from the figure, aside from the ability to automatically start specified IOCs upon booting, a currently unique feature of s6-epics is the ability to let both caproto-based IOCs and EPICS IOCs exit gracefully. With qioc_s6, the management of IOCs controlled by s6-epics can be done through the CA protocol; PyDM OPIs (Figure 7) have also been developed for it, available from the supplementary materials.
Figure 6: A usage example for s6-epics.
5 “Sequencer” IOCs with QueueIOC
In EPICS, “sequencers” based on the seq module are used to implement modules like optics, sscan etc; meanwhile, as has been noted in Section 1, it can be easy to make the state transitions in “sequencers” too complex – comparable to the “goto hell” in C/C++. From our perspective, the most important reason for this is the inexpressiveness of C/C++, the languages EPICS is based on; with QueueIOC the implementation can be much cleaner, by leveraging the expressiveness of Python. We would also like to note that as has been hinted in Figure 4(b), a less noticed aspect of “sequencers” is the automated manipulation of PVs in other IOCs through the CA protocol; in our humble opinion this, instead of state transitions, is the real essence of EPICS “sequencers”. Consequently, in QueueIOC although there is a class QSequencerIOC for general state machines, all “sequencers” discussed in this section are actually based on its subclass QMotorSeqIOC. The latter creates state machine with only two states, “down” and “up”, in order to implement a workalike of seq’s “all channels connected & received 1st monitor”. As the name implies, it is mainly used to implement “sequencers” IOCs that (just like those from optics and sscan) manipulate motor IOCs, which we think is a most common application scenario for them. Similar to their EPICS counterparts, all these IOCs expose some motor-like interface, which we have intentionally designed to be unified; PyDM OPIs (Figure 7, available from the supplementary materials) are given for them, along with a Bluesky encapsulation (available as QueueMotor in the open-source edition of Mamba, cf. Section 4).
Figure 7: PyDM OPIs for the qioc_s6 IOC and QMotorSeqIOC-based IOCs, used in attitude tuning of a simulation of the Raman spectrometer at B5 of HEPS.
The simplest “sequencer” IOCs are used to systematically prevent collision (bumping) between motors; they are based on the class QFuncbumperIOCBase, with qbumper_sim (Figure 8) as a simple example IOC. A more complex QFuncbumperIOCBase-based IOC has been deployed at the transmission X-ray microscope beamline (BE) of HEPS, while more similar IOCs are expected to be deployed at quite a few of the 15 beamlines of HEPS Phase I. As can be seen from qbumper_sim, the complicated logic in the main loop of these “sequencers” IOCs are abstracted in the library (without involving code reminiscent of the “goto hell”), so that developers only need to specify the essential information; in the case of anti-bumping, the information is the mathematical constraints and the list of motor PVs. Under the hood, QFuncbumperIOCBase maintains the state machine in accordance with the underlying motor IOCs; based on this, it denies motion requests when the specified constraints would be violated, when any motor correlated with the requested motor (including itself) is already moving, and when the machine is in the “down” state. Depending on the requirements for other QueueIOC-based “sequencer” IOCs, the information that needs to be specified by the developer can be less mathematical; among them are IOCs based on the class QMotorHubIOCBase, used to control motors connected to multiplexer PLCs (“motor hubs”) for motion controllers. An example IOC based on the class is qmhub_b5, which is actually used in attitude tuning (Li et al., 2024) of the Raman spectrometer at the hard X-ray high-resolution spectroscopy beamline (B5) of HEPS, with a simulated variant in the open-source edition of Mamba (cf. also Figure 67). A particularly notable feature of QMotorHubIOCBase is a delay between the latest motion of a motor and the switch from it to any other motor connected to the same multiplexer as this motor, along with another delay after this switch before the newly chosen motor can be moved; both delays are tunable, and are enforced to prevent potential power surges that may damage the devices involved.
Figure 8: Annotated source code for the qbumper_sim IOC.
Systematic support in QueueIOC for monochromators, a perhaps most widely known application scenario for “sequencer” IOCs, is provided in the class QMonochromatorIOCBase. Based on it, workalikes of the optics IOCs for double-crystal monochromators and high-resolution monochromators are given as qmono_dcm and qmono_hr. For training purposes, a simple “monochromator” IOC is provided as qmono_sim, which involves two axes and simply uses their sum or difference as the “energy” value; a usage example for it is given in the supplementary materials. Apart from the basic features, QMonochromatorIOCBase also provides support for automated speed tuning to make the beam change smoothly when the energy value is changed; of course, just like its counterpart from the optics IOC, this class only implements a linear approximation to the ideal behavior. In qmono_dcm and qmono_hr, support is also provided for the user-friendly specification of Miller indices and lattice parameters. This feature, as well as the speed-tuning feature above, are also provided in ways we believe to be most friendly for IOC developers: eg. the latter is enabled by default, and does not need any code to be written by the developer; it can also be opted out by simply overriding the _auto_velo member of QMonochromatorIOCBase. As can be seen from the source code of these IOCs, in our eyes the essence of monochromators IOCs are coordinate transformations between the energy value and the motor positions; calibration is the process of setting up parameters for the transformations before the latter are put into actual use. Therefore QMonochromatorIOCBase can also be used for multiple-to-multiple coordinate transformations: eg. those between the reciprocal and direct spaces in crystallography, where the transformation parameters may be imported from specialised programs like xrayutilities (https://github.com/dkriegner/xrayutilities) and diffcalc (https://github.com/dls-controls/diffcalc). By exposing transformation parameters as PVs, like what is done in the QMonochromatorIOCBase-based IOCs discussed above, the importing process can be automated. Using the PV interface, the parameters can also be exported to other programs: eg. trajectory programs used in fly scans involving monochromators or the reciprocal space, which can achieve vastly more accurate motion behaviours than what is possible with the speed-tuning feature of QMonochromatorIOCBase.
6 Conclusion
Architectural deficiencies in EPICS lead to inefficiency in development and application; from the perspective of complexity and succinctness, the essence of these problems is that the architecture of EPICS makes the complexities of EPICS IOCs vastly higher than the lower bounds. A backward-compatible way to avoid these problems is replacing EPICS IOCs with Python IOCs based on libraries like caproto. Learning from EPICS OPIs, we can require widgets in GUI frontends to communicate only with the main event loop, and mandate the use of message passing for this communication; based on this idea, the submit/notify pattern is formed, which are also related to the actor/CSP models and the MVC pattern. Mamba frontends and standalone GUI frontends following the pattern have been developed; the communication between both kinds of frontends and their backends may be categorised into requests/replies and notifications, which are handled in main event loops inside the backends. The combination of an event loop, requests and notifications can also be observed elsewhere; thus by treating caput, caget and camonitor as specialised requests/replies and notifications, and handling them in Python-based main loops, the QueueIOC framework is formed. After comparing the architecture of QueueIOC with that of EPICS, we believe QueueIOC has the potential to eventually replace most EPICS IOCs currently used with succinct workalikes; under certain conditions, it may even become a starting point for a unified device-control ecosystem. Examples given for QueueIOC include workalikes of StreamDevice/asyn, as well as seq-like IOCs for monochromators, motor anti-bumping and motor multiplexing; also reported are a workalike of procServ, as well as a procServControl workalike based on it and QueueIOC. A QueueIOC-based framework for detector integration, which overcomes areaDetector’s limitations in performance and architecture, is presented in (Zhang et al., 2024).
Acknowledgements
Liu would like to thank the Linux Club of Peking University for providing a venue in 2019 to share his preliminary thoughts on replacing EPICS (https://lcpu.club/wiki/index.php?title=2019%E6%B4%BB%E5%8A%A8B07). Much more importantly, LCPU introduced him to the wonderland of Unix and Lisp, where the inspirations for this paper came from; finding solutions much simpler than what have been previously imagined, just like those in (Sussman and Steele, 1998) and (Landley, 2017), has become a continuous source of joy and motivation. This work was supported by the National Key Research and Development Program for Young Scientists (Grant No. 2023YFA1609900) and the Young Scientists Fund of the National Natural Science Foundation of China (Grants Nos. 12205328, 12305371).

### v3

@@
 An unintrusive solution is replacing EPICS IOCs with more maintainable and flexible Python IOCs, only reusing the Channel Access (CA) protocol of EPICS.
-After a digression about GUI development inspired by EPICS operator interfaces (OPIs), the structural similarity between standalone GUI backends, the Mamba backend, EPICS IOCs and other server programs is analysed.
-By combining the caproto library and event loops like those in these server programs, the QueueIOC framework for Python IOCs is created, which has the potential to systematically replace most EPICS IOCs currently used.
+After a digression about GUI development inspired by EPICS operator interfaces (OPIs), the structural similarity between standalone GUI backends, the Mamba backend, EPICS IOCs and other server-like programs is analysed.
+By combining the caproto library and event loops like those in these programs, the QueueIOC framework for Python IOCs is created, which has the potential to systematically replace most EPICS IOCs currently used.
 Examples are first given for workalikes of StreamDevice and asyn; examples for seq-like applications include monochromators, motor anti-bumping and motor multiplexing.
 Also shown is software to use with the ~/iocBoot convention which addresses some issues with a similar solution based on procServ, along with a workalike of procServControl.
-A QueueIOC-based framework for detector integration, which overcomes areaDetector’s limitations in performance and architecture, is presented in an accompanying paper.
+A QueueIOC-based framework for detector integration, which aims to overcome some architectural limitations of areaDetector while still offering decent performance, is presented in an accompanying paper.
 1 Introduction
@@
 Figure 4b), which lead to inefficiency in development and application.
-Our first complaints with EPICS are about its conception of “record links” (Kraimer et al., 2018), as well as the locking, scanning and processing mechanism associated with them.
-EPICS record links have different types and attributes: there are input links, output links and forward links; they can be subdivided into database links, Channel Access links and other subtypes; the “Process Passive” attribute can have the values NPP, PP, CA, CP and CPP.
-Events (hardware interrupts, user writing, periodic “scanning”) result in record “processing”, where the selection and ordering of records to be processed are determined by these links, according to the latter’s types and attributes.
-The rules for this selection and ordering are complex, and a major factor contributing to the complexity is the tight coupling between the types and attributes of record links.
-Ideally there should be separate rules about types, subtypes and attributes (especially “Process Passive” among the last) that add up to form the rules for individual combinations; the actual rules are quite far from able to be decomposed (the results would surely be easier to remember), and are instead quite like a lot of special cases piled together.
-The complexity of record links leads to considerable difficulty in learning: even among experienced EPICS application developers, few can precisely retell the rule for a less common combination of type and attribute.
-Fortunately, in practice only a handful of combinations are commonly used; on the other hand, even with the already complex set of rules, there are still requirements that cannot be naturally done, and a prime example for this is the motor module.
+Our first complaints with EPICS are about its conception of “record links” (Kraimer et al., 2018).
+EPICS record links have different types (input links, output links and forward links), subtypes (database links, Channel Access links etc) and attributes (eg.
+NPP, PP, CA, CP and CPP for the “Process Passive” attributes).
+Events (hardware interrupts, user writing, periodic “scanning”) result in record “processing”, where the selection and ordering of records to be processed are determined by these links according to a large and complex rule set.
+The rule set cannot be decomposed into separate rules about types, subtypes and attributes, which leads to considerable difficulty in learning.
+And even with the already complex rule set, there are still requirements that cannot be naturally done, and a prime example for this is the motor module.
 In addition to the types and attributes of links, the selection and ordering of records to be processed are also affected by the records themselves, depending on them being “synchronous” or “asynchronous”; however, motor records are neither synchronous nor asynchronous.
-Apart from the link mechanism, EPICS records themselves are also quite inexpressive: EPICS “databases”, which are collections of records, can be nested; however, nested database files cannot be loaded at runtime, and instead must be expanded at compile time; .substitutions files can be loaded at runtime, but it cannot be nested either.
-This makes it very hard to abstract reusable functionalities as pluggable sub-databases, even for simple requirements like proportional-integral-derivative (PID) controllers.
-One idiomatic alternative is writing these functionalities as customised records, eg.
+Apart from the link mechanism, EPICS records themselves are also quite inexpressive: EPICS “databases”, which are collections of records, cannot be nested at runtime, which makes it very hard to abstract reusable functionalities as pluggable sub-databases, even for simple requirements like proportional-integral-derivative (PID) controllers.
+One idiomatic alternative is to implement customised records, eg.
 epid from the std module, which involves a considerable amount of boilerplate code and quite an extent of knowledge on EPICS internals.
-Another common alternative is writing “sequencers” based on the seq module, which is often error-prone because these “sequencers” can easily become comparable to C/C++ code riddled with goto statements; the monochromator code in the optics module are representative examples for this issue.
+Another common alternative is writing “sequencers” based on the seq module, which is often error-prone (cf.
+Section 6).
 The st.cmd language in EPICS does not have looping or conditional constructs, so we also cannot circumvent the inexpressiveness of records on the st.cmd layer.
-To work around the inexpressiveness of st.cmd itself, it is common practice to use external code generators, eg.
+A common practice to work around this is to use external code generators, eg.
 iocbuilder (Abbott and Cobb, 2011), makeDb.py (from the ADGenICam module) and our own ones (cf.
 ADXspress3 in the next paragraph).
-We note that as can be seen from the examples given, code generators are in particular commonly used to deal with hardware that expose dynamic interfaces: variable numbers of detector channels, variable sets of camera features, etc.
-st.cmd also feels inconvenient to operators, as it is mainly intended as a developer-oriented language, and its inexpressiveness further impedes user-oriented abstraction.
-It is undoubted that none of the issues above is fatal to EPICS: developers can still produce usable, reliable EPICS modules, and users can still use these modules to actually fulfill their needs; as has been summarised above, the problems are about inexpressiveness and inconvenience.
+st.cmd also feels inconvenient to operators, further impeding user-oriented abstraction.
+It is undoubted that none of these issues is fatal to EPICS: developers can still produce usable, reliable EPICS modules, and users can still use these modules to actually fulfill their needs; as has been summarised above, the problems are about inexpressiveness and inconvenience.
 Meanwhile, during the construction of HEPS, we have found that improvements on expressiveness and convenience can be of great value in boosting efficiency; here we briefly give a few examples for this (Figure 1).
 The ihep-pkg packaging framework (Liu, Dong and Li, 2022, cf.
-also Section 4) now covers a large range of EPICS modules than that of the NSLS-II epicsdeb repository (Brookhaven National Laboratory, 2018), and provides support for CentOS 7, Rocky 8 plus partial support for Windows (using the MinGW environment), while still being easy to keep up to date.
+also Section 4) now covers a larger range of EPICS modules than that of the NSLS-II epicsdeb repository (Brookhaven National Laboratory, 2018); it provides support for CentOS 7, Rocky 8 and hopefully compatible environments like RHEL 7/8, as well as partial support for Windows (using the MinGW environment); meanwhile it is still easy to keep up to date, even in comparison with recent works like installSynApps (Derbenev et al., 2023).
 The ADXspress3 module (https://codeberg.org/CasperVector/ADXspress3) merges functionalities from both of its upstream versions (https://github.com/quantumdetectors/xspress3-epics and https://github.com/epics-modules/xspress3), and is much easier to maintain than both upstreams; meanwhile, to users it is also more convenient to build and use.
@@
 Karabo (Göries et al., 2023); for facilities with more legacies, a more backward-compatible option is to reuse the Channel Access (CA) protocol of EPICS, but internally use more succinct tools.
-Given the common programming practices in last decade (especially those related to large scientific facilities), Python is an obvious candidate for this task; in particular, we have chosen the pure-Python caproto library (Allan, 2021), as it provides succinct interfaces for both server-side and client-side programming with the CA protocol.
-However, the officially expected usage of caproto depends on Python’s async/await programming paradigm, which leads to additional costs in learning and development; in this paper, we describe how we solve this problem, and produce a general-purpose workalike of EPICS based on Python.
+Given the common programming practices in last decade (especially those related to large scientific facilities), Python is an obvious candidate for this task; in particular, we have chosen the pure-Python caproto library (Allan, 2021), as it provides succinct interfaces for both server-side and client-side programming with the CA protocol; alternatives like PCASPy (https://github.com/paulscherrerinstitute/pcaspy) may have issues like https://github.com/pyepics/pyepics/issues/176.
+However, the officially expected usage of caproto depends on Python’s async/await programming paradigm, which leads to additional costs in learning and development, although it is very helpful in highly concurrent application scenarios; in this paper, we describe how we solve this, and produce a general-purpose workalike of EPICS based on Python.
 2 The submit/notify pattern for GUI programming
@@
 The “controller” part is the submit/notify logic separated into the widgets’ event callbacks (or slot functions in the signal/slot mechanism) and the main loop.
-As is hinted above, the main event loop in a GUI frontend should only contain logic and states essential to the graphical representation of the business logic – which in turn belongs to the backend; this improves the reliability and maintainability of GUIs, just like how privilege separation improves the security of qmail and software with similar designs (Bernstein, 2007).
+As is hinted above, the main event loop in a GUI frontend should only contain logic and states essential to the graphical representation of the business logic – which in turn belongs to the backend; this improves the reliability and maintainability of GUIs.
 In the Mamba framework, the backend communicates with the frontends through a remote-procedure call (RPC) mechanism (Figure 2a) based on ZeroMQ.
@@
 Section 4); released in the same repository are two PyQt-based examples for standalone submit/notify GUIs.
-One of them is a workalike of the program in (Du et al., 2012), essentially an X-ray beam position monitor (XBPM) based on area detectors, with a simplified RPC mechanism between the backend and the frontend based on Python’s queue.
-The another is a simplified workalike of the ImageJ plugin provided by the ADViewers module, treating EPICS IOCs for area detectors as its backends, reusing the CA protocol for message passing with them.
+One of them is a workalike of the program in (Du et al., 2012), essentially an X-ray beam-position monitor (XBPM) based on area detectors, with a simplified RPC mechanism between the backend and the frontend based on Python’s queue.
+The other is a simplified workalike of the ImageJ plugin provided by the ADViewers module, treating areaDetector IOCs as its backends, reusing the CA protocol for message passing with them.
 3 The client-server model and EPICS IOCs
 As has been discussed in Section 2, the CA protocol can be seen as a message-passing mechanism between servers (EPICS IOCs) and clients (EPICS OPIs etc); moreover, IOCs may be compared to GUI backends (whether standalone or in Mamba), and sometimes even be directly used as the backends.
-Therefore from analysing what is common in these servers, it is possible to find better ways to do what EPICS IOCs do; historically our standalone backends also learned from the Mamba backend, so here we begin with Mamba.
+Therefore from analysing what is common in these server-like programs, it is possible to find better ways to do what EPICS IOCs do; historically our standalone backends also learned from the Mamba backend, so here we begin with Mamba.
 Communication between the backend and frontends in Mamba may be categorised into requests/replies and notifications, the latter necessary due to the intrinsic weaknesses of state polling (Liu et al., 2022).
@@
 We find the combination of an event loop, requests (with or without replies; an example for the latter is submissions in the submit/notify pattern, cf.
-Section 2) and notifications a common pattern in server programs.
+Section 2) and notifications a common pattern in server-like programs.
 In addition to GUI backends, other examples are also abundant, eg.
@@
 Figure 3: The alsamixer program.
-In the light of these server programs, it is easy to find that caput, caget and camonitor, the basic operations in the CA protocol of EPICS, are also specialised combinations of requests/replies and notifications.
+In the light of these server-like programs, it is easy to find that caput, caget and camonitor, the basic operations in the CA protocol of EPICS, are also specialised combinations of requests/replies and notifications.
 Given these observations, we designed the basic architecture of our Python IOC framework, as in Figure 4(a).
@@
 Correspondingly, there is a thin layer that isolates from the main loop the async/await-based code caproto expects (cf.
-Section 1); as the layer is based on variants of Python’s queue, we call our framework QueueIOC.
+Section 1); as the layer is based on various message queues, we call our framework QueueIOC.
 For many IOCs in our framework the main loops are explicit, and are quite like the main loop in the standalone XBPM program in Section 2; however, for certain types of requirements exhibiting obvious regularity (eg.
@@
 Figure 1, 5 & 8).
-Figure 4: (a) The architecture of QueueIOC, in comparison with (b) the architecture of an EPICS IOC; the latter image is courtesy of Kraimer et al. (2018).
-Aside from QueueIOC and barebone caproto, we are also aware of similar attempts at using Python in the EPICS ecosystem, eg.
+Figure 4: (a) The architecture of QueueIOC, in comparison with (b) the architecture of an EPICS IOC; the latter image is courtesy of Kraimer et al.
+(2018).
+We note that although (b) is for the stagnant 3.15.x branch of EPICS, its actively developed 7.x branch is not fundamentally different in terms of architecture (except for the introduction of the PVA protocol aside from CA), and is also affected by the issues summarised in this paper.
+Aside from QueueIOC, (barebone) caproto and PCASPy, we are also aware of other attempts at using Python in the EPICS ecosystem, eg.
 PyDevice (Vodopivec, 2020) and pythonSoftIOC (Cobb, 2021).
-Instead of comparing them in detail, here we note the most crucial difference between QueueIOC and the rest is that it attempts to replace most EPICS IOCs in a systematic way, while striving to keep the amounts of code for the Python IOCs satisfactorily close to their intuitive complexity lower-bounds.
+Instead of comparing them in detail, here we note the most crucial difference between QueueIOC and the rest is that it attempts to replace most EPICS IOCs (in the narrow sense, structured like what makeBaseApp.pl creates) in a systematic way, while striving to keep the amounts of code for the Python IOCs satisfactorily close to their intuitive complexity lower-bounds.
 For the latter goal, we explicitly note that QueueIOC focuses on fulfilling needs in succinct ways, which are not necessarily how EPICS fulfills the same needs.
 In the following sections, concrete examples will be given for various types of applications doable with QueueIOC; we believe these examples can adequately show the potential of QueueIOC.
+Here we explicitly note that what we want is not to aggressively replace all existing EPICS IOCs, but instead to provide a smooth transition path to more efficient alternatives that are (protocol-wise) compatible.
 We also note that motor IOCs are perhaps a most obvious type of IOCs currently not covered by QueueIOC, but that we do not find this a fundamental weakness.
@@
 With mechanisms like QScanIOC (cf.
-Section 4) and QSlowIOC (Zhang et al., 2024), what scanners do can also be done succinctly with QueueIOC.
+Section 4) and QSlowIOC (Zhang et al., 2024), what “scanners” do can also be done succinctly with QueueIOC.
 In summary, all architectural elements in EPICS IOCs have satisfactory counterparts in QueueIOC; so in the long term, we believe QueueIOC has the potential to be eventually capable of doing what is done with all EPICS IOCs currently used, except those IOCs that run under special resource constraints (most prominently VME IOCs).
-Furthermore, as has been noted above, caput, caget and camonitor are just specialised requests/replies and notifications; we think similar conclusions could be made about the PV Access (PVA) protocol added to EPICS in recent years.
-In conjunction with our observations about server programs, it is also a reasonable guess that similar conclusions could be made about the communication protocols in other device-control ecosystems, eg.
+There may also be performance-related concerns about QueueIOC due to its Python-based nature, but we find it unnecessary to worry about: with the IOC qscan_rate available from the supplementary materials, it can be verified that QueueIOC can easily offer refresh (“scan”) rates up to at least 100 Hz and monitor rates up to at least 500 Hz, both of which are comparable to what EPICS IOCs are normally expected to do; with QDetectorIOC (Section 4), QueueIOC also offers decent detector readout performance comparable to areaDetector.
+Our outlook can also be extended beyond the CA protocol of EPICS: as has been noted above, caput, caget and camonitor are just specialised requests/replies and notifications; we think similar conclusions could be made about the PV Access (PVA) protocol added to EPICS in recent years.
+In conjunction with our observations about server-like programs, it is also a reasonable guess that similar conclusions could be made about the communication protocols in other device-control ecosystems, eg.
 TANGO and Karabo.
@@
 with cmd_map in qscan_b2985, which would be clumsy to do with EPICS databases.
+A more complex example is the qscan_hfda IOC, which will be detailed in Section 6.
 Based on the extensive support for communication protocols available with Python (whether in standard libraries or third-party libraries), with QueueIOC it is also easy to write IOCs based on more complex interfaces like JSON, HTTP, modbus etc: eg.
 the qdet_eiger IOC for the Eiger detector uses the vendor’s protocols based on JSON/HTTP and ZeroMQ.
-qdet_eiger is based on QDetectorIOC, a QueueIOC-based framework for detector integration presented in an accompanying paper (Zhang et al., 2024), which overcomes the areaDetector framework’s limitations in performance and architecture.
-Also demonstrated therein is how QueueIOC works with C/C++ libraries from vendors.
+qdet_eiger is based on QDetectorIOC, a QueueIOC-based framework for detector integration presented in an accompanying paper (Zhang et al., 2024), which aims to overcome some architectural limitations of the areaDetector framework while still offering decent performance.
 Figure 5: Source code for the qscan_b2985 IOC.
@@
 QueueIOC’s automatic adaption to dynamic hardware interfaces (cf.
-also Figure 8) is based on Python’s type() mechanism for runtime creation of classes; based on this mechanism, we can do advanced abstractions similar to those in the ADGenICam and ADEiger IOCs with much more succinct code (Zhang et al., 2024).
+also Figure 8) is based on Python’s type() mechanism for runtime creation of classes; based on this mechanism, we can also create advanced abstractions for variable sets of detector features, similar to those in the ADGenICam and ADEiger IOCs but with much more succinct code (Zhang et al., 2024).
 As can be seen from Figure 5, IOCs based on QueueIOC, whether for devices with dynamic interfaces or not, are easy to run for operators: unlike st.cmd in EPICS, with these Python IOCs device identifiers (addresses, ports etc) and other tuning parameters can be given on the command line, which is well separated from the IOCs’ source code.
@@
 5 “Sequencer” IOCs with QueueIOC
-In EPICS, “sequencers” based on the seq module are used to implement modules like optics, sscan etc; meanwhile, as has been noted in Section 1, it can be easy to make the state transitions in “sequencers” too complex – comparable to the “goto hell” in C/C++.
-From our perspective, the most important reason for this is the inexpressiveness of C/C++, the languages EPICS is based on; with QueueIOC the implementation can be much cleaner, by leveraging the expressiveness of Python.
+In EPICS, “sequencers” based on the seq module are used to implement modules like optics, sscan etc; meanwhile, as will be detailed in Section 6, writing seq-based “sequencers” can be error-prone.
+With QueueIOC the implementation can be much cleaner, by leveraging the expressiveness of Python and following a succinct approach.
 We would also like to note that as has been hinted in Figure 4(b), a less noticed aspect of “sequencers” is the automated manipulation of PVs in other IOCs through the CA protocol; in our humble opinion this, instead of state transitions, is the real essence of EPICS “sequencers”.
@@
 The latter creates state machine with only two states, “down” and “up”, in order to implement a workalike of seq’s “all channels connected & received 1st monitor”.
-As the name implies, it is mainly used to implement “sequencers” IOCs that (just like those from optics and sscan) manipulate motor IOCs, which we think is a most common application scenario for them.
+As the name implies, it is mainly used to implement “sequencers” IOCs that (just like those in optics and sscan) manipulate motor IOCs, which we think is a most common application scenario for them.
 Similar to their EPICS counterparts, all these IOCs expose some motor-like interface, which we have intentionally designed to be unified; PyDM OPIs (Figure 7, available from the supplementary materials) are given for them, along with a Bluesky encapsulation (available as QueueMotor in the open-source edition of Mamba, cf.
@@
 A more complex QFuncbumperIOCBase-based IOC has been deployed at the transmission X-ray microscope beamline (BE) of HEPS, while more similar IOCs are expected to be deployed at quite a few of the 15 beamlines of HEPS Phase I.
-As can be seen from qbumper_sim, the complicated logic in the main loop of these “sequencers” IOCs are abstracted in the library (without involving code reminiscent of the “goto hell”), so that developers only need to specify the essential information; in the case of anti-bumping, the information is the mathematical constraints and the list of motor PVs.
+As can be seen from qbumper_sim, the complicated logic in the main loop of these “sequencers” IOCs are abstracted in the library, so that developers only need to specify the essential information; in the case of anti-bumping, the information is the mathematical constraints and the list of motor PVs.
 Under the hood, QFuncbumperIOCBase maintains the state machine in accordance with the underlying motor IOCs; based on this, it denies motion requests when the specified constraints would be violated, when any motor correlated with the requested motor (including itself) is already moving, and when the machine is in the “down” state.
@@
 Systematic support in QueueIOC for monochromators, a perhaps most widely known application scenario for “sequencer” IOCs, is provided in the class QMonochromatorIOCBase.
-Based on it, workalikes of the optics IOCs for double-crystal monochromators and high-resolution monochromators are given as qmono_dcm and qmono_hr.
-For training purposes, a simple “monochromator” IOC is provided as qmono_sim, which involves two axes and simply uses their sum or difference as the “energy” value; a usage example for it is given in the supplementary materials.
-Apart from the basic features, QMonochromatorIOCBase also provides support for automated speed tuning to make the beam change smoothly when the energy value is changed; of course, just like its counterpart from the optics IOC, this class only implements a linear approximation to the ideal behavior.
+Based on it, workalikes of the optics IOCs for double-crystal monochromators and high-resolution monochromators are respectively given as qmono_dcm and qmono_hr; a usage example for the latter is given in the supplementary materials.
+For training purposes, a simple “monochromator” IOC is provided as qmono_sim, which involves two axes and simply uses their sum or difference as the “energy” value.
+Apart from the basic features, QMonochromatorIOCBase also provides support for automated speed tuning to make the beam change smoothly when the energy value is changed; of course, just like its counterpart in the optics IOC, this class only implements a linear approximation to the ideal behavior.
 In qmono_dcm and qmono_hr, support is also provided for the user-friendly specification of Miller indices and lattice parameters.
@@
 trajectory programs used in fly scans involving monochromators or the reciprocal space, which can achieve vastly more accurate motion behaviours than what is possible with the speed-tuning feature of QMonochromatorIOCBase.
+6 Case studies: some laser controllers and monochromators
+In Section 3, we noted that QueueIOC attempts to replace most EPICS IOCs in a systematic way, while striving to keep the new IOCs as simple as reasonable.
+We realise that after the brief tour of functionalities in Sections 45, it may still not be obvious what the unique benefits of QueueIOC are; so in this section, we analyse these benefits in detail by having a close look at some example IOCs.
+Our first example is the qscan_hfda IOC for CNI PSU-H-FDA and PSU-A-D laser controllers.
+In this IOC, the ancillary class PyHfda and function make_hfda() (Figure 9) are more notable than the IOC class HfdaIOCBase.
+PSU-H-FDA (older) and PSU-A-D (newer) share a basic serial-based communication protocol, but the older model does not support state readback (cf.
+the function read_state() in PyHfda); the state readback provides more than 10 information items, all of which are mapped into PVs by the IOC.
+The state-changing commands require checksum bytes and produce echo replies, so in case a setpoint item is not available from the state readback, the IOC needs to decide whether to update the value of the corresponding PV upon user writing based on the correctness of the replies to the commands.
+For this kind of readback-less setpoints (enable_val, as well as current_val and power_val for the older model), the IOC also needs to properly initialise the corresponding hardware settings, so that the initial values of their PVs are correctly reflected.
+Moreover, the newer model is known to encounter read timeouts if a state-reading command is sent too closely after a state-changing command, so an additional delay should be enforced after the latter.
+All the above are handled succinctly and cleanly in qscan_hfda: state readback in read_state() based on a call to struct.unpack(); complete correctness check in send_cmd() and read_state() based on checksums and echo replies; the automatic differentiation between the old and new models, as well as the preparation of initial states, in make_hfda().
+Done in HfdaIOCBase are the automatic construction of PV lists for different models, the different handling of state changing depending on the model, and the delay after state changing.
+While these are not impossible in EPICS IOCs, they would be very awkward to implement: what the struct.unpack() does above would be highly bloated with StreamDevice; additionally StreamDevice cannot easily support the echo check and the delay after state changing, and we also do not find seq-based “sequencers” very helpful on this issue.
+Writing the IOC in C/C++ may be also considered, but the codebase would still be significantly larger, especially considering the support for different versions; these make the development and maintenance cost of such an IOC much higher than our IOC.
+Figure 9: Some notable fragments of the qscan_hfda IOC.
+What distinguishes QueueIOC from the others is an intentional pursuit of the utmost simplicity (Hoare, 1981), or in other words approaching the complexity lower-bounds (cf.
+Section 1); what QueueIOC attempts to do, in practice, is to provide succinct yet powerful tools that developers can use to build IOCs that are close to their complexity lower-bounds.
+Admittedly, with caproto, PCASPy etc, it would not be hard to write workalikes of qscan_hfda that are not much more complex than it.
+This is because its logic revolves around simple state-changing and state-reading commands, and the abstraction for this pattern, QScanIOC, is easy to reimplement; as has been hinted above, PyHfda and make_hfda() are what really make the IOC simple.
+There are also examples where the tools provided by QueueIOC significantly contribute to the simplicity of the IOCs: in fact, all IOCs for monochromators, motor anti-bumping and motor multiplexing in Section 5 can be seen as examples for this; it can be easily seen when the reader considers how the same requirements could be implemented with EPICS, caproto, PCASPy etc.
+Here we also discuss monochromators in detail, as another practical demonstration of how simplicity is achieved in QueueIOC-based IOCs – with not only particular tools, but also reusable design patterns.
+Writing seq-based “sequencers” is often error-prone, because the state transitions in them can easily become comparable to the “goto hell” C/C++ (Figure 10c), result in hard-to-trace bugs (a few of them are given in the supplementary materials).
+This is why most “sequencers” in QueueIOC only have the “down” and “up” states (cf.
+Section 5); the domain-specific business logic is instead encapsulated in the state-transition functions.
+In monochromator IOCs based on QMonochromatorIOCBase, the functions serve_down(), serve_up(), do_common() etc are the basis of our abstraction of their business logic as coordinate transformations.
+However, in addition to these shared logic, real-world monochromators need invertible three-way conversions between the Bragg angle $\theta$, the wavelength $\lambda$ and the energy $E$; the conversions can be even more complex for certain monochromators, eg.
+the high-resolution monochromator (Figure 10a).
+To do this cleanly, we use functions to encode these conversions: eg.
+mono_phis() in the qmono_hr IOC (Figure 10b), as well as mono_theta() which is used by qmono_dcm and usable by other simple variants of the double-crystal monochromator.
+Figure 10: (a) Conversion between the wavelength $\lambda$, the energy $E$, the Bragg angles $\theta_{1}$, $\theta_{2}$ and the motorised angles $\varphi_{1}$, $\varphi_{2}$ of the high-resolution monochromator in its “rock $\theta_{2}$” mode.
+(b) Example usage of mono_phis(), where the 2nd argument specifies a symmetric monochromator geometry, while the 3rd and 4th arguments respectively specify the Bragg spacings $2d_{1}$ and $2d_{2}$.
+(c) State transitions of the hrCtl “sequencer” in optics.
+In comparison with the monochromator IOCs in optics, IOCs like qmono_dcm are not only much shorter, but also much better modularised.
+The shared logic is contained in the code of QMonochromatorIOCBase, organised cleanly unlike the “goto hell”.
+The $\theta$/$\lambda$/$E$ conversions are fully encapsulated in functions like mono_phis(), which are implemented as straightforward encodings of graphs like Figure 10(a).
+The rest cannot be easily abstracted as libraries, but they are still organised cleanly, with a structure intentionally kept consistent across different monochromator IOCs; this can be seen from a comparison between the IOC classes, eg.
+MonoHrIOC in qmono_hr and MonoDcmIOC in qmono_dcm.
+The result of the structural differences above is dramatic reduction in the cost of development and maintenance.
+When writing a new monochromator IOC, the developer often only needs to customise the conversion function and the IOC class, without worrying about the possibility of a small change conflicting with some assumptions implicit in the global state.
+When debugging, issues can usually be easily traced and resolved thanks to the careful decoupling: eg.
+the conversion function can isolated and used as a calculator for the variables involved, greatly facilitating tests.
+The same approach is also followed elsewhere in QueueIOC for non-trivial IOCs: aside from PyHfda in qscan_hfda, another example is QDetectorIOC (Zhang et al., 2024), where C/C++ libraries (conventionally called “SDKs” or software development kits) from vendors are encapsulated thinly into self-made mini-“SDKs”; the latter expose succinct, reusable and adaptable interfaces, which can be tested in isolation and even reused in standalone applications.
+With measures taken like those above, the IOCs provided by QueueIOC have been made satisfactorily close to their complexity lower-bounds.
+As has been noted in Section 1, improvements in simplicity coincide with improvements in efficiency; with the consistent pursuit of simplicity throughout QueueIOC, we believe it is able to bring about significant efficiency boosts in the EPICS ecosystem.
 7 Conclusion
@@
 A backward-compatible way to avoid these problems is replacing EPICS IOCs with Python IOCs based on libraries like caproto.
-Learning from EPICS OPIs, we can require widgets in GUI frontends to communicate only with the main event loop, and mandate the use of message passing for this communication; based on this idea, the submit/notify pattern is formed, which are also related to the actor/CSP models and the MVC pattern.
+Learning from EPICS OPIs, we can require widgets in GUI frontends to communicate only with the main event loop, and mandate the use of message passing for this communication; based on this idea, the submit/notify pattern is formed, which is also related to the actor/CSP models and the MVC pattern.
 Mamba frontends and standalone GUI frontends following the pattern have been developed; the communication between both kinds of frontends and their backends may be categorised into requests/replies and notifications, which are handled in main event loops inside the backends.
@@
 Examples given for QueueIOC include workalikes of StreamDevice/asyn, as well as seq-like IOCs for monochromators, motor anti-bumping and motor multiplexing; also reported are a workalike of procServ, as well as a procServControl workalike based on it and QueueIOC.
-A QueueIOC-based framework for detector integration, which overcomes areaDetector’s limitations in performance and architecture, is presented in (Zhang et al., 2024).
+A QueueIOC-based framework for detector integration, which aims to overcome some architectural limitations of areaDetector while still offering decent performance, is presented in (Zhang et al., 2024).
+A practical analysis is given for the unique benefits of QueueIOC, emphasising its pursuit of the utmost simplicity, which leads to significant reduction in the cost of development and maintenance.
 Acknowledgements

### v4

@@
 Abstract
-Architectural deficiencies in EPICS lead to inefficiency in the development and application of EPICS input/output controllers (IOCs).
-An unintrusive solution is replacing EPICS IOCs with more maintainable and flexible Python IOCs, only reusing the Channel Access (CA) protocol of EPICS.
-After a digression about GUI development inspired by EPICS operator interfaces (OPIs), the structural similarity between standalone GUI backends, the Mamba backend, EPICS IOCs and other server-like programs is analysed.
-By combining the caproto library and event loops like those in these programs, the QueueIOC framework for Python IOCs is created, which has the potential to systematically replace most EPICS IOCs currently used.
-Examples are first given for workalikes of StreamDevice and asyn; examples for seq-like applications include monochromators, motor anti-bumping and motor multiplexing.
-Also shown is software to use with the ~/iocBoot convention which addresses some issues with a similar solution based on procServ, along with a workalike of procServControl.
-A QueueIOC-based framework for detector integration, which aims to overcome some architectural limitations of areaDetector while still offering decent performance, is presented in an accompanying paper.
+Keywords:
+EPICS; Python IOC; submit/notify pattern; s6-epics; software architecture.
+Background and Purpose:
+Architectural deficiencies in EPICS lead to inefficiency in the development and application of EPICS IOCs.
+An unintrusive solution is replacing EPICS IOCs with more maintainable and flexible Python IOCs, only reusing the CA protocol of EPICS.
+While there are libraries like caproto and PCASPy that help to create Python IOCs, they still feel insufficient for more complex requirements.
+Methods:
+Noticing caput, caget and camonitor are just specialised combinations of requests/replies and notifications in client-server communication, by combining barebone caproto and event loops like those in server-like programs, the QueueIOC framework for Python IOCs is created, which has the potential to systematically reduce the development and maintenance cost of IOCs.
+Results:
+Examples based on QueueIOC are first given for workalikes of StreamDevice and asyn; also given are examples for “sequencer” applications, like those based on seq, include monochromators, motor anti-bumping and motor multiplexing.
+A QueueIOC-based framework for detector integration is presented in an accompanying paper.
+Also reported is a simple but expressive architecture for GUIs, as well as software to use with the ~/iocBoot convention which addresses some issues we find with a similar solution based on procServ.
 1 Introduction
-The Experimental Physics and Industrial Control System (EPICS) is a basis for accelerator control and beamline control at HEPS, the High Energy Photon Source (Chu et al., 2018; Liu, Dong and Li, 2022).
+The Experimental Physics and Industrial Control System (EPICS) is a basis for accelerator control and beamline control at HEPS, the High Energy Photon Source [1, 2].
 The rich selection of EPICS modules available has greatly facilitated device-control tasks at HEPS.
@@
 Figure 4b), which lead to inefficiency in development and application.
-Our first complaints with EPICS are about its conception of “record links” (Kraimer et al., 2018).
+Our first complaints with EPICS are about its conception of “record links” [3].
 EPICS record links have different types (input links, output links and forward links), subtypes (database links, Channel Access links etc) and attributes (eg.
-NPP, PP, CA, CP and CPP for the “Process Passive” attributes).
+NPP, PP, CA, CP and CPP for the “Process Passive” attribute).
 Events (hardware interrupts, user writing, periodic “scanning”) result in record “processing”, where the selection and ordering of records to be processed are determined by these links according to a large and complex rule set.
@@
 One idiomatic alternative is to implement customised records, eg.
-epid from the std module, which involves a considerable amount of boilerplate code and quite an extent of knowledge on EPICS internals.
+epid from the std module, which involves a considerable amount of repeated code and quite an extent of knowledge on EPICS internals.
 Another common alternative is writing “sequencers” based on the seq module, which is often error-prone (cf.
@@
 A common practice to work around this is to use external code generators, eg.
-iocbuilder (Abbott and Cobb, 2011), makeDb.py (from the ADGenICam module) and our own ones (cf.
+iocbuilder [4], makeDb.py (from the ADGenICam module) and our own ones (cf.
 ADXspress3 in the next paragraph).
@@
 Meanwhile, during the construction of HEPS, we have found that improvements on expressiveness and convenience can be of great value in boosting efficiency; here we briefly give a few examples for this (Figure 1).
-The ihep-pkg packaging framework (Liu, Dong and Li, 2022, cf.
-also Section 4) now covers a larger range of EPICS modules than that of the NSLS-II epicsdeb repository (Brookhaven National Laboratory, 2018); it provides support for CentOS 7, Rocky 8 and hopefully compatible environments like RHEL 7/8, as well as partial support for Windows (using the MinGW environment); meanwhile it is still easy to keep up to date, even in comparison with recent works like installSynApps (Derbenev et al., 2023).
+The ihep-pkg packaging framework [2] now covers a larger range of EPICS modules than that of the NSLS-II epicsdeb repository [5]; meanwhile it is still easy to keep up to date, even in comparison with recent works like installSynApps [6].
+It provides support for CentOS 7, Rocky 8 and hopefully compatible environments like RHEL 7/8; partial support is provided for Windows, Debian/Ubuntu and Rocky/RHEL 9, mainly targeting scenarios where the C/C++ libraries (conventionally called “SDKs” or software development kits) from vendors mandate the use of these environments.
 The ADXspress3 module (https://codeberg.org/CasperVector/ADXspress3) merges functionalities from both of its upstream versions (https://github.com/quantumdetectors/xspress3-epics and https://github.com/epics-modules/xspress3), and is much easier to maintain than both upstreams; meanwhile, to users it is also more convenient to build and use.
-The MambaPlanner mechanism (Li et al., 2023) provides a succinct, consistent command-line interface (CLI) for both step scans and fly scans, encapsulating hairy details about hardware configuration, correctness checks and data processing; while being convenient for debugging and advanced usage, it also facilitates development of graphical user interfaces (GUIs) oriented toward regular users.
-As is shown in Figure 1, they have each approached some kind of complexity lower-bound, in the sense that there is no obvious way to significantly simplify the code without inordinately complicating other code modules or sacrificing code clarity.
-Following the “succinctness is power” statement (Graham, 2002), by approaching the complexity lower-bound we can maximise the engineering efficiency.
-The examples above are just natural results of this rule; the efficiency boosts in them can be in 1–2 orders of magnitude, which also nicely coincide with the reduction of code needed to do the same tasks.
+The MambaPlanner mechanism [7] provides a succinct, consistent command-line interface (CLI) for both step scans and fly scans, encapsulating hairy details about hardware configuration, correctness checks and data processing; while being convenient for debugging and advanced usage, it also facilitates development of graphical user interfaces (GUIs) oriented toward regular users.
+As is shown in Figure 1, they have each approached some kind of complexity/cost lower-bound, in the sense that there is no obvious way to significantly simplify the code without inordinately complicating other code modules or sacrificing code clarity.
+In the examples above, the efficiency boosts in them can be in 1–2 orders of magnitude, which also nicely coincide with the reduction of code needed to do the same tasks.
 Similarly, in other applications at HEPS, improvements in succinctness also prove to coincide with improvements in efficiency, even when the complexity lower-bounds have not yet been approached.
+More profoundly, as has been discussed recently [8], we believe that the pursuit to minimise complexity/cost is a pursuit of the collaboration and coevolution of artificial intelligence (AI) with human intelligence.
 Figure 1: Some code examples, with the essential information in bold.
 From the perspective above, we can give a summary of the problems with EPICS: the essence of these problems is that its architecture makes the complexities of EPICS modules vastly higher than the lower bounds; the latter can often be estimated intuitively, just like in Figure 1.
-We understand that the design of EPICS is deeply affected by historical factors: the choice of record and links over a full-fledged language, as well as the inexpressiveness of st.cmd, may be largely seen as results of hardware limitations on Versa Module Europa (VME) and similar hardware platforms; record and links are also reminiscent of the programming paradigm common with programmable logical controllers (PLCs), PandABox (Christian et al., 2019) etc.
+We understand that the design of EPICS is deeply affected by historical factors: the choice of record and links over a full-fledged language, as well as the inexpressiveness of st.cmd, may be largely seen as results of hardware limitations on Versa Module Europa (VME) and similar hardware platforms; record and links are also reminiscent of the programming paradigm common with programmable logical controllers (PLCs), PandABox [9] etc.
 Nevertheless, with the fast growth of EPICS usage on PCs, it is also time to rethink about these design decisions.
 Of course, an option for completely new facilities is to embrace ecosystems with less historical burdens, eg.
-Karabo (Göries et al., 2023); for facilities with more legacies, a more backward-compatible option is to reuse the Channel Access (CA) protocol of EPICS, but internally use more succinct tools.
-Given the common programming practices in last decade (especially those related to large scientific facilities), Python is an obvious candidate for this task; in particular, we have chosen the pure-Python caproto library (Allan, 2021), as it provides succinct interfaces for both server-side and client-side programming with the CA protocol; alternatives like PCASPy (https://github.com/paulscherrerinstitute/pcaspy) may have issues like https://github.com/pyepics/pyepics/issues/176.
+Karabo [10]; for facilities with more legacies, a more backward-compatible option is to reuse the Channel Access (CA) protocol of EPICS, but internally use more succinct tools.
+Given the common programming practices in last decade (especially those related to large scientific facilities), Python is an obvious candidate for this task; in particular, we have chosen the pure-Python caproto library [11], as it provides succinct interfaces for both server-side and client-side programming with the CA protocol; alternatives like PCASPy (https://github.com/paulscherrerinstitute/pcaspy) may have issues like https://github.com/pyepics/pyepics/issues/176.
 However, the officially expected usage of caproto depends on Python’s async/await programming paradigm, which leads to additional costs in learning and development, although it is very helpful in highly concurrent application scenarios; in this paper, we describe how we solve this, and produce a general-purpose workalike of EPICS based on Python.
@@
 In this section, we will discuss a GUI programming pattern inspired by EPICS, which in turn helped to inspire our Python-based EPICS workalike besides being useful in itself.
-In our summary, the complexity of GUI programming has a few common sources, the first of which is the mixing of business logic with operations on widgets that represent the logic.
+In our summary, the complexity in GUI programming has a few common sources, the first of which is the mixing of business logic with operations on widgets that represent the logic.
 This kind of mix-up is a quite well-known type of non-modularity in software engineering; it is generally resolved by decoupling GUI programs into frontends and backends, eg.
-those in Mamba (Liu et al., 2022; Dong et al., 2022).
+those in Mamba [12, 13].
 Another source of complexity is the complex interactions between widgets in GUI frontends, which complicates the call chains between these widgets.
@@
 GUIs in the EPICS ecosystem, known as operator interfaces (OPIs), can be easily composed in OPI editors in a drag-and-drop fashion.
-As drag and drop are inefficient in the creation of large OPIs with many repetitions, we are also systematically exploring PyDM (Slepicka, 2018), the Python-based OPI engine for this task.
+As drag and drop are inefficient in the creation of large OPIs with many repetitions, we are also systematically exploring PyDM [14], the Python-based OPI engine for this task.
 Notwithstanding this issue, we still find GUIs an advantage of EPICS, since we do not need to worry about the problems in the previous paragraph: OPI widgets do not directly interact or share states with each other, eliminating problems about call chains and timing.
@@
 Our attempts to systematically simplify the control logic in IOCs, by means of caproto-based Python IOCs, will be covered in the following sections, and here we continue our discussion on GUIs.
-Learning from OPIs, even in a general GUI we can require the widgets to communicate only with the main event loop (Figure 2a); to minimise state sharing, message passing can be mandated for this communication.
+Learning from OPIs, even in a general GUI we can require widgets to communicate only with the main event loop (Figure 2a); to minimise state sharing, message passing can be mandated for this communication.
 To perform message passing, libraries like ZeroMQ and Python’s queue can surely be used; the signal/slot mechanism provided by some GUI libraries, eg.
@@
 For tasks (eg.
-some in beamline control) that do not need the full Mamba infrastructure or even barebone Bluesky (Allan et al., 2019), we also developed some standalone GUIs with the submit/notify pattern; in these GUIs, there is still a clear separation between frontends and backends.
-Mamba frontends developed in the submit/notify style are, as of now, primarily used in our framework for automated attitude tuning of beamlines, presented in an accompanying paper (Li et al., 2024).
-The code for these frontends has been released in the open-source edition of Mamba (cf.
-Section 4); released in the same repository are two PyQt-based examples for standalone submit/notify GUIs.
-One of them is a workalike of the program in (Du et al., 2012), essentially an X-ray beam-position monitor (XBPM) based on area detectors, with a simplified RPC mechanism between the backend and the frontend based on Python’s queue.
+some in beamline control) that do not need the full Mamba infrastructure or even barebone Bluesky [15], we also developed some standalone GUIs with the submit/notify pattern; in these GUIs, there is still a clear separation between frontends and backends.
+Mamba frontends developed in the submit/notify style are, as of now, primarily used in our framework for automated attitude tuning of beamlines [8].
+The code for these frontends has been released in the open-source edition of Mamba; released in the same repository are two PyQt-based examples for standalone submit/notify GUIs.
+One of them is a workalike of the program in [16], essentially an X-ray beam-position monitor (XBPM) based on area detectors, with a simplified RPC mechanism between the backend and the frontend based on Python’s queue.
 The other is a simplified workalike of the ImageJ plugin provided by the ADViewers module, treating areaDetector IOCs as its backends, reusing the CA protocol for message passing with them.
@@
 Therefore from analysing what is common in these server-like programs, it is possible to find better ways to do what EPICS IOCs do; historically our standalone backends also learned from the Mamba backend, so here we begin with Mamba.
-Communication between the backend and frontends in Mamba may be categorised into requests/replies and notifications, the latter necessary due to the intrinsic weaknesses of state polling (Liu et al., 2022).
+Communication between the backend and frontends in Mamba may be categorised into requests/replies and notifications, the latter necessary due to the intrinsic weaknesses of state polling [12].
 In both standalone and Mamba backends there are main event loops: in Mamba backends, event handling (also including sending notifications) is separated into handlers registered to the core library, which encapsulates the main event loop; in standalone backends, there are explicit main loops that reply to requests and send notifications.
@@
 For many IOCs in our framework the main loops are explicit, and are quite like the main loop in the standalone XBPM program in Section 2; however, for certain types of requirements exhibiting obvious regularity (eg.
-the QScanIOC-based IOCs in Section 4 and all IOCs in Section 5), we also encapsulate the boilerplate code in their main loops, so that the developer usually only needs to care about the essentials (cf.
+the QScanIOC-based IOCs in Section 4 and all IOCs in Section 5), we also abstract and encapsulate the repeated code in their main loops, so that the developer usually only needs to care about the essentials (cf.
 Figure 1, 5 & 8).
-Figure 4: (a) The architecture of QueueIOC, in comparison with (b) the architecture of an EPICS IOC; the latter image is courtesy of Kraimer et al.
-(2018).
-We note that although (b) is for the stagnant 3.15.x branch of EPICS, its actively developed 7.x branch is not fundamentally different in terms of architecture (except for the introduction of the PVA protocol aside from CA), and is also affected by the issues summarised in this paper.
+Figure 4: (a) The architecture of QueueIOC, in comparison with (b) the architecture of an EPICS IOC; the latter image is courtesy of [3].
+Although (b) is for the stagnant 3.15.x branch of EPICS, we do not find its actively developed 7.x branch fundamentally different in terms of architecture (except for the introduction of the PVA protocol aside from CA); the 7.x branch is also affected by the issues summarised in this paper.
 Aside from QueueIOC, (barebone) caproto and PCASPy, we are also aware of other attempts at using Python in the EPICS ecosystem, eg.
-PyDevice (Vodopivec, 2020) and pythonSoftIOC (Cobb, 2021).
+PyDevice [17] and pythonSoftIOC [18].
 Instead of comparing them in detail, here we note the most crucial difference between QueueIOC and the rest is that it attempts to replace most EPICS IOCs (in the narrow sense, structured like what makeBaseApp.pl creates) in a systematic way, while striving to keep the amounts of code for the Python IOCs satisfactorily close to their intuitive complexity lower-bounds.
@@
 With mechanisms like QScanIOC (cf.
-Section 4) and QSlowIOC (Zhang et al., 2024), what “scanners” do can also be done succinctly with QueueIOC.
+Section 4) and QSlowIOC [19], what “scanners” do can also be done succinctly with QueueIOC.
 In summary, all architectural elements in EPICS IOCs have satisfactory counterparts in QueueIOC; so in the long term, we believe QueueIOC has the potential to be eventually capable of doing what is done with all EPICS IOCs currently used, except those IOCs that run under special resource constraints (most prominently VME IOCs).
@@
 TANGO and Karabo.
-If these guesses were sufficiently accurate, QueueIOC might become a starting point for some kind of unified device-control ecosystem, since QueueIOC intentionally encourages a programming style agnostic of the device-control ecosystem: the developer usually only needs to care about requests/replies and notifications – not caput, caget or camonitor.
+If these guesses were sufficiently accurate, QueueIOC might become a starting point for some kind of unified device-control ecosystem, since QueueIOC intentionally encourages a programming style agnostic of the device-control ecosystem: the developer usually only needs to care about requests/replies and notifications – not caput, caget or camonitor (cf.
+also the discussion on modularity in Section 6).
 4 Device IOCs and “soft” IOCs with QueueIOC
-QueueIOC and s6-epics (the latter to be covered below) have been released, respectively, at https://codeberg.org/CasperVector/queue_iocs and https://codeberg.org/CasperVector/s6-epics; fully open-source editions of Mamba and ihep-pkg have been released, respectively, at https://codeberg.org/CasperVector/mamba-ose and https://codeberg.org/CasperVector/ihep-pkg-ose.
-QueueIOC itself depends on some patches for caproto; GUIs in the open-source edition of Mamba (including the standalone ones, which are in the mamba_lite subdirectory) depend on some patches for pyqtgraph; the supplementary materials include OPIs which depend on some patches for PyDM; all these currently HEPS-specific patches are available from the open-source edition of ihep-pkg.
 Our actual introduction to QueueIOC begins with a workalike of StreamDevice, which is in our eyes a simplest yet most useful EPICS module: the QScanIOC class, in combination with classes like TimedRWPair; an example IOC for this, qscan_b2985, is given for the Keysight B2985 electrometer, shown in Figure 5.
@@
 Unlike StreamDevice which uses “protocol files”, essentially a still quite limiting domain-specific language (DSL), QScanIOC-based IOCs have native access to Python’s capability to process textual and/or binary data.
-As these IOCs also have full access to Python’s expressiveness, we can easily abstract boilerplate code, eg.
+As these IOCs also have full access to Python’s expressiveness, we can easily abstract repeated code, eg.
 with cmd_map in qscan_b2985, which would be clumsy to do with EPICS databases.
@@
 the qdet_eiger IOC for the Eiger detector uses the vendor’s protocols based on JSON/HTTP and ZeroMQ.
-qdet_eiger is based on QDetectorIOC, a QueueIOC-based framework for detector integration presented in an accompanying paper (Zhang et al., 2024), which aims to overcome some architectural limitations of the areaDetector framework while still offering decent performance.
+qdet_eiger is based on QDetectorIOC, a QueueIOC-based framework for detector integration presented in an accompanying paper [19], which aims to overcome some architectural limitations of the areaDetector framework while still offering decent performance.
 Figure 5: Source code for the qscan_b2985 IOC.
@@
 This is why the main loop in QScanIOC does not need to be explicitly written for each device; for most other devices, explicit main loops are necessary.
-For training purposes in writing main loops, the qdet_nct IOC for the Tsuji (N)CT counter and the qdet_o974 IOC for the Ortec 974 counter are good examples, both based on QDetectorIOC.
+For training purposes in writing main loops, the qseq_bestec IOC in Section 6 is an expressive example; the qdet_nct IOC for the Tsuji (N)CT counter and the qdet_o974 IOC for the Ortec 974 counter are also good examples, both based on QDetectorIOC.
 QScanIOC is not used as we want them to support areaDetector-like acquisition of a fixed (and possibly infinite) number of frames, a feature unfit for stateless implementations.
@@
 Section 1).
-QueueIOC’s automatic adaption to dynamic hardware interfaces (cf.
-also Figure 8) is based on Python’s type() mechanism for runtime creation of classes; based on this mechanism, we can also create advanced abstractions for variable sets of detector features, similar to those in the ADGenICam and ADEiger IOCs but with much more succinct code (Zhang et al., 2024).
-As can be seen from Figure 5, IOCs based on QueueIOC, whether for devices with dynamic interfaces or not, are easy to run for operators: unlike st.cmd in EPICS, with these Python IOCs device identifiers (addresses, ports etc) and other tuning parameters can be given on the command line, which is well separated from the IOCs’ source code.
+QueueIOC’s automatic adaption to variable device interfaces (cf.
+also Figure 8, the qseq_bestec IOC and the self-made SDKs in Section 6) is based on Python’s mechanisms for runtime creation and using of classes; based on these mechanisms, we can also create advanced abstractions for variable sets of detector features, similar to those in IOCs based on ADGenICam but with much more succinct code [19].
+As can be seen from Figure 5, IOCs based on QueueIOC, whether for devices with variable interfaces or not, are easy to run for operators: unlike st.cmd in EPICS, with these Python IOCs device identifiers (addresses, ports etc) and other tuning parameters can be given on the command line, which is well separated from the IOCs’ source code.
 Additionally, when necessary it is also quite easy to customise IOC behaviours on deeper levels, without needing to copy entire IOC source files: eg.
@@
 Till now all IOCs discussed in this section interact with hardware other than the controlling computer; besides “sequencers” which will be covered in Section 5, pure-software IOCs can also be written with QueueIOC.
-Our example for pure-software QueueIOC-based IOCs is qioc_s6, a workalike for the procServControl IOC, based on s6-epics, a workalike of procServ (Thompson, 2004).
+Our example for pure-software QueueIOC-based IOCs is qioc_s6, a workalike for the procServControl IOC, based on s6-epics, a workalike of procServ [20].
 s6-epics itself is based on s6 (https://skarnet.org/software/s6/), a well-designed suite of programs to manage service processes; that latter most importantly (to us and concerning EPICS) supports separate logging for each service with reliable log rotation to avoid exhaustion of disk space.
-After organising IOCs in the ~/iocBoot convention, we can use administration commands like those in (Liu, Dong and Li, 2022), and a few additional useful commands (Figure 6).
+After organising IOCs in the ~/iocBoot convention, we can use administration commands like those in [2], and a few additional useful commands (Figure 6).
 As can be seen from the figure, aside from the ability to automatically start specified IOCs upon booting, a currently unique feature of s6-epics is the ability to let both caproto-based IOCs and EPICS IOCs exit gracefully.
@@
 We would also like to note that as has been hinted in Figure 4(b), a less noticed aspect of “sequencers” is the automated manipulation of PVs in other IOCs through the CA protocol; in our humble opinion this, instead of state transitions, is the real essence of EPICS “sequencers”.
-Consequently, in QueueIOC although there is a class QSequencerIOC for general state machines, all “sequencers” discussed in this section are actually based on its subclass QMotorSeqIOC.
+Consequently, in QueueIOC although there is a class QSequencerIOC for general state machines (cf.
+also the qseq_bestec IOC in Section 6), all “sequencers” discussed in this section are actually based on its subclass QMotorSeqIOC.
 The latter creates state machine with only two states, “down” and “up”, in order to implement a workalike of seq’s “all channels connected & received 1st monitor”.
 As the name implies, it is mainly used to implement “sequencers” IOCs that (just like those in optics and sscan) manipulate motor IOCs, which we think is a most common application scenario for them.
-Similar to their EPICS counterparts, all these IOCs expose some motor-like interface, which we have intentionally designed to be unified; PyDM OPIs (Figure 7, available from the supplementary materials) are given for them, along with a Bluesky encapsulation (available as QueueMotor in the open-source edition of Mamba, cf.
-Section 4).
+Similar to their EPICS counterparts, all these IOCs expose some motor-like interface, which we have intentionally designed to be unified; PyDM OPIs (Figure 7, available from the supplementary materials) are given for them, along with a Bluesky encapsulation (available as QueueMotor in the open-source edition of Mamba).
 Figure 7: PyDM OPIs for the qioc_s6 IOC and QMotorSeqIOC-based IOCs, used in attitude tuning of a simulation of the Raman spectrometer at B5 of HEPS.
@@
 Depending on the requirements for other QueueIOC-based “sequencer” IOCs, the information that needs to be specified by the developer can be less mathematical; among them are IOCs based on the class QMotorHubIOCBase, used to control motors connected to multiplexer PLCs (“motor hubs”) for motion controllers.
-An example IOC based on the class is qmhub_b5, which is actually used in attitude tuning (Li et al., 2024) of the Raman spectrometer at the hard X-ray high-resolution spectroscopy beamline (B5) of HEPS, with a simulated variant in the open-source edition of Mamba (cf.
+An example IOC based on the class is qmhub_b5, which is actually used in attitude tuning [8] of the Raman spectrometer at the hard X-ray high-resolution spectroscopy beamline (B5) of HEPS, with a simulated variant in the open-source edition of Mamba (cf.
 also Figure 67).
@@
 Using the PV interface, the parameters can also be exported to other programs: eg.
-trajectory programs used in fly scans involving monochromators or the reciprocal space, which can achieve vastly more accurate motion behaviours than what is possible with the speed-tuning feature of QMonochromatorIOCBase.
-6 Case studies: some laser controllers and monochromators
+trajectory programs used by motion controllers in variable-speed fly scans involving monochromators or the reciprocal space, which can achieve vastly more accurate motion behaviours than what is possible with the speed-tuning feature of QMonochromatorIOCBase.
+6 Case studies: some controllers and “sequencers”
 In Section 3, we noted that QueueIOC attempts to replace most EPICS IOCs in a systematic way, while striving to keep the new IOCs as simple as reasonable.
 We realise that after the brief tour of functionalities in Sections 45, it may still not be obvious what the unique benefits of QueueIOC are; so in this section, we analyse these benefits in detail by having a close look at some example IOCs.
-Our first example is the qscan_hfda IOC for CNI PSU-H-FDA and PSU-A-D laser controllers.
+Our first example is the qscan_hfda IOC for CNI PSU-H-FDA, PSU-H-LED and PSU-A-D laser controllers.
 In this IOC, the ancillary class PyHfda and function make_hfda() (Figure 9) are more notable than the IOC class HfdaIOCBase.
-PSU-H-FDA (older) and PSU-A-D (newer) share a basic serial-based communication protocol, but the older model does not support state readback (cf.
-the function read_state() in PyHfda); the state readback provides more than 10 information items, all of which are mapped into PVs by the IOC.
-The state-changing commands require checksum bytes and produce echo replies, so in case a setpoint item is not available from the state readback, the IOC needs to decide whether to update the value of the corresponding PV upon user writing based on the correctness of the replies to the commands.
-For this kind of readback-less setpoints (enable_val, as well as current_val and power_val for the older model), the IOC also needs to properly initialise the corresponding hardware settings, so that the initial values of their PVs are correctly reflected.
-Moreover, the newer model is known to encounter read timeouts if a state-reading command is sent too closely after a state-changing command, so an additional delay should be enforced after the latter.
-All the above are handled succinctly and cleanly in qscan_hfda: state readback in read_state() based on a call to struct.unpack(); complete correctness check in send_cmd() and read_state() based on checksums and echo replies; the automatic differentiation between the old and new models, as well as the preparation of initial states, in make_hfda().
+While the 3 models share a basic serial-based communication protocol, only PSU-A-D supports state readback, and there are also minor differences in the ways to set the laser intensity; the state readback (cf.
+the function read_state() in PyHfda) provides more than 10 information items, all of which are mapped into PVs by the IOC.
+The state-changing commands require checksum bytes and produce echo replies (the models also differ slightly in their echo behaviours, as can be seen from send_cmd() in PyHfda), so in case a setpoint item is not available from the state readback, the IOC needs to decide whether to update the value of the corresponding PV upon user writing based on the correctness of the replies to the commands.
+For this kind of readback-less setpoints (enable_val, as well as current_val and power_val for models other than PSU-A-D), the IOC also needs to properly initialise the corresponding hardware settings, so that the initial values of their PVs are correctly reflected.
+Moreover, the controllers may encounter read timeouts if a state-reading command is sent too closely after a state-changing command, so an additional delay should be enforced after the latter.
+All the above are handled succinctly and cleanly in qscan_hfda: state readback in read_state() based on a call to struct.unpack(); complete correctness check in send_cmd() and read_state() based on checksums and echo replies; the automatic differentiation between the models, as well as the preparation of initial states, in make_hfda().
 Done in HfdaIOCBase are the automatic construction of PV lists for different models, the different handling of state changing depending on the model, and the delay after state changing.
@@
 Figure 9: Some notable fragments of the qscan_hfda IOC.
-What distinguishes QueueIOC from the others is an intentional pursuit of the utmost simplicity (Hoare, 1981), or in other words approaching the complexity lower-bounds (cf.
-Section 1); what QueueIOC attempts to do, in practice, is to provide succinct yet powerful tools that developers can use to build IOCs that are close to their complexity lower-bounds.
 Admittedly, with caproto, PCASPy etc, it would not be hard to write workalikes of qscan_hfda that are not much more complex than it.
 This is because its logic revolves around simple state-changing and state-reading commands, and the abstraction for this pattern, QScanIOC, is easy to reimplement; as has been hinted above, PyHfda and make_hfda() are what really make the IOC simple.
-There are also examples where the tools provided by QueueIOC significantly contribute to the simplicity of the IOCs: in fact, all IOCs for monochromators, motor anti-bumping and motor multiplexing in Section 5 can be seen as examples for this; it can be easily seen when the reader considers how the same requirements could be implemented with EPICS, caproto, PCASPy etc.
+There are also examples where the tools provided by QueueIOC significantly contribute to the simplicity of the IOCs: in fact, all IOCs in Section 5 can be seen as examples for this; it can be easily seen when the reader considers how the same requirements could be implemented with EPICS, caproto, PCASPy etc.
+Another example is the “sequencer” IOC qseq_bestec available from the supplementary materials, used with a Bestec optical complex (PGM, slit & mirror unit) at the high-resolution nanoscale electronic structure spectroscopy beamline (BC) of HEPS.
+The vendor supplies a proprietary IOC-like program that encapsulate the device interfaces as waveform PVs, which are essentially argument lists for control commands.
+As what we really want are scalar PVs, this IOC was written to translate the argument lists in “master” PVs into single elements in “slave” PVs, and translate caput operations on slave PVs back into caput on the master PVs.
+In this IOC, a table of master-PVs and their slave-PVs is first created; then accordingly, the slave PVs are created and the master PVs are connected to; in the main event loop, caput operation on the slave PVs and monitored updates of the master PVs are handled.
+This IOC is simple and easily adaptable to another set of master and slave PVs; this is because all the caput and monitor events are represented in a well-designed way there, allowing for a single function to handle all the PVs without repeated code.
+The similar requirements would be much less convenient to implement, whether with caproto, PCASPy etc or (especially) with EPICS and seq.
+In our summary, what distinguishes QueueIOC from the others is an intentional and systematic pursuit of the utmost simplicity [21], or in other words approaching the complexity/cost lower-bounds (cf.
+Section 1); what QueueIOC attempts to do, in practice, is to provide succinct yet powerful tools that developers can use to build IOCs that are close to their complexity/cost lower-bounds.
 Here we also discuss monochromators in detail, as another practical demonstration of how simplicity is achieved in QueueIOC-based IOCs – with not only particular tools, but also reusable design patterns.
@@
 The shared logic is contained in the code of QMonochromatorIOCBase, organised cleanly unlike the “goto hell”.
-The $\theta$/$\lambda$/$E$ conversions are fully encapsulated in functions like mono_phis(), which are implemented as straightforward encodings of graphs like Figure 10(a).
+The $\theta$$\lambda$/$E$ conversions are fully encapsulated in functions like mono_phis(), which are implemented as straightforward encodings of graphs like Figure 10(a).
 The rest cannot be easily abstracted as libraries, but they are still organised cleanly, with a structure intentionally kept consistent across different monochromator IOCs; this can be seen from a comparison between the IOC classes, eg.
@@
 the conversion function can isolated and used as a calculator for the variables involved, greatly facilitating tests.
-The same approach is also followed elsewhere in QueueIOC for non-trivial IOCs: aside from PyHfda in qscan_hfda, another example is QDetectorIOC (Zhang et al., 2024), where C/C++ libraries (conventionally called “SDKs” or software development kits) from vendors are encapsulated thinly into self-made mini-“SDKs”; the latter expose succinct, reusable and adaptable interfaces, which can be tested in isolation and even reused in standalone applications.
-With measures taken like those above, the IOCs provided by QueueIOC have been made satisfactorily close to their complexity lower-bounds.
-As has been noted in Section 1, improvements in simplicity coincide with improvements in efficiency; with the consistent pursuit of simplicity throughout QueueIOC, we believe it is able to bring about significant efficiency boosts in the EPICS ecosystem.
+The same approach is also followed elsewhere in QueueIOC for non-trivial IOCs: aside from PyHfda in qscan_hfda, another example is the qscan_redpower IOC for SPI redPOWER laser controllers based on helpers/qdev_redpower_lib.py, a self-made mini-“SDK”.
+In the complex feature set provided by the controllers, this SDK covers a non-trivial subset currently demanded by the high-pressure beamline (B6) of HEPS, where those controllers are used.
+Similarly, in QDetectorIOC [19], vendors’ SDKs are also encapsulated thinly into mini-SDKs; many of these SDKs can dynamically adapt to variable device interfaces.
+These self-made SDKs expose succinct, reusable and adaptable interfaces, which can be tested in isolation and even reused in standalone applications.
+In summary, modularity facilitates reuse, which enhances regularity and therefore helps to reduce the complexity of the entire system (cf.
+also the discussion on PVA, TANGO and Karabo in Section 3).
+With measures taken like those above, the IOCs provided by QueueIOC have been made satisfactorily close to their complexity/cost lower-bounds.
+As has been discussed in Section 1, with the consistent pursuit of simplicity throughout QueueIOC, we believe it is able to bring about significant efficiency boosts in the EPICS ecosystem, and help large scientific facilities (along with their staffs) become more intelligent.
 7 Conclusion
@@
 After comparing the architecture of QueueIOC with that of EPICS, we believe QueueIOC has the potential to eventually replace most EPICS IOCs currently used with succinct workalikes; under certain conditions, it may even become a starting point for a unified device-control ecosystem.
-Examples given for QueueIOC include workalikes of StreamDevice/asyn, as well as seq-like IOCs for monochromators, motor anti-bumping and motor multiplexing; also reported are a workalike of procServ, as well as a procServControl workalike based on it and QueueIOC.
-A QueueIOC-based framework for detector integration, which aims to overcome some architectural limitations of areaDetector while still offering decent performance, is presented in (Zhang et al., 2024).
+Examples given for QueueIOC include workalikes of StreamDevice/asyn, as well as “sequencer” IOCs (like those based on seq), including those for monochromators, motor anti-bumping and motor multiplexing; also reported are a workalike of procServ, as well as a procServControl workalike based on it and QueueIOC.
+A QueueIOC-based framework for detector integration, which aims to overcome some architectural limitations of areaDetector while still offering decent performance, is presented in [19].
 A practical analysis is given for the unique benefits of QueueIOC, emphasising its pursuit of the utmost simplicity, which leads to significant reduction in the cost of development and maintenance.
-Acknowledgements
+Statements and declarations
+Acknowledgements:
 Liu would like to thank the Linux Club of Peking University for providing a venue in 2019 to share his preliminary thoughts on replacing EPICS (https://lcpu.club/wiki/index.php?title=2019%E6%B4%BB%E5%8A%A8B07).
-Much more importantly, LCPU introduced him to the wonderland of Unix and Lisp, where the inspirations for this paper came from; finding solutions much simpler than what have been previously imagined, just like those in (Sussman and Steele, 1998) and (Landley, 2017), has become a continuous source of joy and motivation.
+Much more importantly, LCPU introduced him to the wonderland of Unix and Lisp, where the inspirations for this paper came from; finding solutions much simpler than what have been previously imagined, just like those in [22] and [23], has become a continuous source of joy and motivation.
+Funding:
 This work was supported by the National Key Research and Development Program for Young Scientists (Grant No.
@@
 12205328, 12305371).
+Data availability:
+QueueIOC and s6-epics have been released, respectively, at https://codeberg.org/CasperVector/queue_iocs and https://codeberg.org/CasperVector/s6-epics; fully open-source editions of Mamba and ihep-pkg have been released, respectively, at https://codeberg.org/CasperVector/mamba-ose and https://codeberg.org/CasperVector/ihep-pkg-ose.
+QueueIOC itself depends on some patches for caproto; GUIs in the open-source edition of Mamba (including the standalone ones, which are in the mamba_lite subdirectory) depend on some patches for pyqtgraph; the supplementary materials include OPIs which depend on some patches for PyDM; all these currently HEPS-specific patches are available from the open-source edition of ihep-pkg.